OpenStudy (anonymous):

I am doing a review sheet can someone help me with this question? Given f(x) and g(x), write g(f(x)). f(x)=3x-4; g(x)= square root of x+2

OpenStudy (jhannybean):

\[g(f(x)) = g(3x-4) = \sqrt{3x-4}+2\]

OpenStudy (jhannybean):

Wherever you see an \(x\) in the function of g(x), you replace it with the function of f.

OpenStudy (anonymous):

Oh! Well that is a lot easier than I thought! Thank you so much!

OpenStudy (jhannybean):

No problem ^^